# Repeating partial disruptions and two-body relaxation

### v1

@@ 2.2 Two-body relaxation and loss cone theory
The dynamical evolution and relaxation of a nuclear cluster due to two-body encounters is simply and accurately described by the orbit-averaged Fokker-Planck equation (Cohn & Kulsrud(1978)). By directly computing the expected stochastic variations of a test star’s orbital parameters, it can be shown that relaxation in squared angular momentum (normalized to the value $J^{2}_{c}$ at the circular orbit of energy $E$) $\mathcal{R}\equiv J^{2}/J_{c}^{2}$ is more efficient than relaxation in $E$ for eccentric orbits. One can identify two time scales (Merritt(2013)):
@@
where $D=D(E)$ is the orbit averaged diffusion coefficient, that for very eccentric orbits characterises the effects of 2-body encounters (Merritt(2013))
@@
to compute the rate of tidal disruptions at each energy. The relation between $\mathcal{R}_{0}$ and $\mathcal{R}_{\mathrm{TDE}}$ is directly determined by the model of instantaneous tidal disruptions at the pericentre. The rates are computed from the solution $f(E)$ of the 1D evolution in energy (Vasiliev(2017); Stone et al.(2020)Stone, Vasiliev, Kesden, Rossi, Perets, & Amaro-Seoane)
@@
We summarize here the key assumptions made to derive the relaxed profile in Cohn & Kulsrud(1978):
@@ 2. CK 2
Figure 2: The value of the loss cone occupation fraction $q$ as a function of energy (in units of the velocity dispersion at the influence radius) in a system with a central massive black hole of mass $M_{\bullet}=4\times 10^{6}M_{\odot}$ and for a sun-like star. The upper axis shows the value of the radius of the circular orbit at energy $E$. The details of the underlying distributions of stars are in the text. We mark with dotted lines from left to right: in blue the energy corresponding to a circular orbit at the influence radius, in orange the energy where the flux entering the loss cone is maximum, and in green the energy where $q=1$.
@@
We will now consider a prototypical system to compute the strength of 2-body relaxation. The system is composed of a central MBH with mass $M_{\bullet}=4\times 10^{6}M_{\odot}$ surrounded by a distribution of stars (total number $8\times 10^{7}$, individual mass $1M_{\odot}$) and stellar black holes (total number $8\times 10^{4}$, individual mass $10M_{\odot}$) both following a Dehnen profile (Dehnen(1993); Tremaine et al.(1994)Tremaine, Richstone, Byun, Dressler, Faber, Grillmair, Kormendy, & Lauer) with scale radius of $8.8\ \mathrm{pc}$ and inner slope $1.5$, as in Broggi et al.(2022)Broggi, Bortolas, Bonetti, Sesana, & Dotti. This system has an influence radius [^5] There are multiple definitions of influence radius in literature. In this work we define $r_{\mathrm{inf}}=G\,M_{\bullet}/\sigma^{2}_{\mathrm{inf}}$, where $\sigma_{\mathrm{inf}}$ is the local velocity dispersion. of $r_{\mathrm{inf}}\simeq 2.2\ \mathrm{pc}$ and the velocity dispersion at the influence radius corresponds to $\sigma_{\mathrm{inf}}=88\ \mathrm{km/s}$, computed according to the $M-\sigma$ by Gültekin et al.(2009)Gültekin, Richstone, Gebhardt, Lauer, Tremaine, Aller, Bender, Dressler, Faber, Filippenko, Green, Ho, Kormendy, Magorrian, Pinkney, & Siopis. This system is used to compute the quantity $\left\rvert\Delta\mathcal{R}\right\rvert_{orb}$. The loss cone diffusivity $q=q(E)$ in this system is shown in Fig. 2.
@@
In Fig. 3 we show the predictions of our semi-analytical model for partial disruptions of a star. We consider both model A and model B stars; different values of the initial energy $E_{0}$ corresponding to $q=\{0.1,1.0,10.0\}$; different prescriptions for energy change in the core kick (Eqs. 10 and 11); and different models for tidal excitation (linear and non-linear).
@@
Irrespective of the specific models used for the computation of energy changes (i.e. tidal excitation and mass loss), there is a qualitative difference between the partial disruption of stars in the two stellar models – such a difference is driven by the different internal structure. The more compact structure of model $B$ requires deeper penetration of the orbit (smaller $r_{\mathrm{p}}$) for the star to lose mass, so that the function $\delta m/m$ goes to zero at large pericentres faster than to stars described by model A. Moreover, the compact structure of model B corresponds to a larger energy stored into oscillations, with tidal oscillations being dominant at large $r_{\mathrm{p}}$. On the other hand, partial disruptions of model $A$ stars are dominated by the effects of mass loss at larger pericentres. In general, the linear expansion seems to perform well compared to non-linear estimates when mass loss is below approximately $5\%$. While in the case of model A Eq. 11 predicts a larger energy increase with respect to Eq. 10, the opposite is true in the case of model B given the larger pericentres involved in units of $r_{\mathrm{TDE}}$.
@@
To overview the full energy range, in Fig. 4 we show the qualitative effect of single passages at the pericentre as a function of the initial energy $E$ and angular momentum $J$ for model A and model B; we also report the radius $r_{c}$ of a circular orbit at energy $E$ for better interpretation. Here, we use the combination of energy change given by Eq. 11 and non-linear tidal excitations of Eq. 12. We plot the contours that correspond to $\Delta E/E=-100\%$, $|\Delta E/E|=10\%$ and $\delta m/m<1\%$, and define some (arbitrary) thresholds to classify the possible outcomes:
@@
  • Negligible change in the orbit, significant change in the mass. This happens when $|\Delta E|/E<10\%$ and $|\delta m|>1\%$ (aqua green).
  • Significant change in the orbit, significant change in the mass. This happens when $10\%<|\Delta E|/E<100\%$ and $|\delta m|>1\%$ (green).
@@
As with the previous analysis, we see that the two stellar models show different qualitative behaviours for large pericentres. In particular, we see that model A stars have smaller energy fluctuations at $\delta m<1\%$, while the tidal excitation is strongly relevant in model B for semimajor axis (estimated through $r_{c}$) comparable to the influence radius. On the other hand, mass loss effects are dominant for model A and the total energy change is positive (i.e, the resulting orbit is more bound) only at small pericentre passages. Close to $r_{\mathrm{p}}/r_{\mathrm{TDE}}=\eta_{\mathrm{TTDE}}$, the energy fluctuation steeply decreases and the remnant is ejected in both cases.
@@ 3.1 Full loss cone regime
When $q\gtrsim 1$, relaxation acts on timescales shorter than times between two partial disruption events. In this regime, single passages grazing or entering the loss cone radius will alter the orbital parameters once per period $P$, while the timescale of relaxation in angular momentum is $P/q$. This implies that only a small fraction of stars with pericentre smaller than the tidal disruption radius will actually reach the turning point, but they are likely to be scattered to a different (safe) orbit before. As a consequence, the probability of having a pTDE and staying on the same orbit to undergo a second disruption is small, and virtually negligible since $q$ grows quickly as the orbital energy $E$ decreases (see Fig. 2).
A fraction of partial disruptions of loosely bound stars may result in the ejection of the remnant at a pericentre larger than $r_{\mathrm{TDE}}$ (see Fig. 4), and this may be compatible with classical loss cone theory.
In fact, since they both imply an instantaneous removal of the subject star from the distribution, ejections can be treated together with total disruptions in the standard approach to the loss cone, where particles are removed when they reach a pericentre below the critical value $r_{\mathrm{TDE}}$. Therefore, by identifying a suitable, larger threshold radius $r_{\mathrm{LC}}>r_{\mathrm{TDE}}$ that roughly distinguishes the region of ejections plus total disruptions, classical loss cone theory may be simply adapted to provide a good description of the system. Such a value of $r_{\mathrm{LC}}$ weakly depends on energy, but this does not affect the estimation of the corresponding Cohn-Kulsrud solution.
@@ 3.2 Empty loss cone regime
Figure 5: Effects of repeated partial disruptions with initial pericentre $r_{\mathrm{p}}$ (in units of $r_{\mathrm{TDE}}$) according to our simple step-by-step evolution for $q=0.1$ (upper figure), $q=0.01$ (middle figure) and $q=0.001$ (lower figure); the corresponding energy for our fiducial model is also reported in units of $\sigma^{2}_{\mathrm{inf}}$ and the orbital period $P$ in years. The first column refers to model A stars, the second column to model B stars. In the upper panel of each figure, we show the outcome of particles completing less than $N_{q}=1/q$ orbits, distinguishing between total disruptions ($r_{\mathrm{p}}<r_{\mathrm{TDE}}$) and ejections ($E_{\mathrm{fin}}<0$); we dub as survival the stars that complete $N_{q}$ orbits. In the middle panel, we report the ratio between the final and initial energy, showing the $90\%$ line (blue) and the $110\%$ line (orange) for reference. In the lower panel, we show the fraction of completed orbits $N_{\mathrm{orb}}/N_{q}$. We mark with a vertical dashed line the threshold pericentre, i.e. the smallest pericentre allowing the stars to complete $N_{q}$ orbits. Model A stars show a reduced orbital energy as they start closer to the threshold pericentre, meaning that they are less bound to the central MBH. On the other hand, model B stars show significant energy fluctuations slightly biased towards less bound orbits when computing the effect of asymmetric mass loss according to Gafton et al.(2015)Gafton, Tejeda, Guillochon, Korobkin, & Rosswog.
@@
For each pericentre passage, we will consider the mass lost according to Eq. (5) (model A) or Eq. (7) (model B) and the two prescriptions for the corresponding decrease in specific binding energy via core kicks, also including the exchange of energy between stellar oscillations and orbital motion according to Eq. (12). We remark that the perturbation due to tidal excitation is positive-definite (i.e. the orbit becomes more bound) only at the first pericentre passage. At subsequent passages, whether the tidal deformation injects or subtracts energy to the stellar orbit depends on the phase of the ongoing oscillation (induced by the previous passage) when the star is gets back to pericentre (Lai(1997)). To include this effect, we use Eq. 12 to determine the amplitude of the energy exchanged but, from the second pericentre passage onwards, we give it a random sign at each passage.
@@
We treat $R_{\star}$ and $\eta_{\mathrm{TTDE}}$ as fixed parameters throughout the evolution, since modelling the readjustment of the star requires to directly solve the complex hydrodynamic evolution of the stellar structure [^6] Since we are considering orbital periods of at most $10^{4}$ yr, much less than the Kelvin-Helmholtz timescale ($\sim 1-100$ Myr; Kippenhahn & Weigert(1990)), dramatic mass loss and energy injection through the dissipation of nonlinear oscillation modes can in principle seriously affect the internal structure of the star.. In Appendix A we explore a non-fiducial scenario where the star can inflate due to nonlinear tides. In this case, part of the energy stored in oscillations is converted into internal energy through an enlargement of the radius. Finally, we remark that this procedure implies that the specific angular momentum is constant throughout the repeated disruptions.
In Fig. 5 we show the results of our simulations by reporting the outcome (survival, total disruption or ejection), the final energy and the number of repeating partial disruption orbits for each simulation as a function of its initial pericentre $r_{\mathrm{p}}$. For both models and for any value of $q$, we see that repeating the partial disruption process can lead to total disruption and, in a minority of cases, to ejection at pericentres larger than $r_{\mathrm{TDE}}$ and $\eta_{\mathrm{TDE}}\,r_{\mathrm{TDE}}$ The threshold radius (reported in units of $r_{\mathrm{TDE}}$ in each sub-figure) increases with $N_{q}$ and therefore decreases with $q$. There are, however, some significant differences between the repeated partial disruption of a star described by model A and one described by model B.

### v2

@@ 2.2 Two-body relaxation and loss cone theory
The dynamical evolution and relaxation of a nuclear cluster due to two-body encounters is simply and accurately described by the orbit-averaged Fokker-Planck equation (Cohn & Kulsrud, 1978). By directly computing the expected stochastic variations of a test star’s orbital parameters, it can be shown that relaxation in squared angular momentum (normalised to the value $J^{2}_{c}$ at the circular orbit of energy $E$) $\mathcal{R}\equiv J^{2}/J_{c}^{2}$ is more efficient than relaxation in $E$ for eccentric orbits. One can identify two time scales (Merritt, 2013):
@@
where $D=D(E)$ is the orbit averaged diffusion coefficient, that for very eccentric orbits characterises the effects of 2-body encounters (Merritt, 2013)
@@
to compute the rate of tidal disruptions at each energy. The relation between $\mathcal{R}_{0}$ and $\mathcal{R}_{\mathrm{TDE}}$ is directly determined by the model of instantaneous tidal disruptions at the pericentre. The rates are computed from the solution $f(E)$ of the 1D evolution in energy (Vasiliev, 2017; Stone et al., 2020)
@@
We summarise here the key assumptions made to derive the relaxed profile in Cohn & Kulsrud (1978):
@@ 2. CK 2
Figure 3: The value of the loss cone diffusivity $q$ as a function of energy (in units of the velocity dispersion at the influence radius) in a system with a central massive black hole of mass $M_{\bullet}=4\times 10^{6}M_{\odot}$ and for a sun-like star. The stellar distribution is composed of Solar mass stars and stellar mass black holes, both following a Dehnen profile with inner slope $\rho\propto r^{-1.5}$. We show the curve $q(E)$ both for the initial conditions (blue) and after $10$ Gyr of relaxation (orange). The upper axis shows the radius of the circular orbit at energy $E$. The details of the underlying distributions of stars are in the text. We mark with vertical dotted lines from left to right: in blue, the energy corresponding to a circular orbit at the influence radius; in orange, the energy where the flux entering the loss cone is maximised; and in green, the energy where $q=1$.
@@
We will now consider a prototypical system to compute the strength of 2-body relaxation. The system is composed of a central MBH with mass $M_{\bullet}=4\times 10^{6}M_{\odot}$ surrounded by a distribution of stars (total number $8\times 10^{7}$, individual mass $1M_{\odot}$) and stellar black holes (total number $8\times 10^{4}$, individual mass $10M_{\odot}$) both following a Dehnen profile (Dehnen, 1993; Tremaine et al., 1994) with scale radius of $8.8\ \mathrm{pc}$ and inner slope $1.5$, as in Broggi et al. (2022). This system has an influence radius  [^5] There are multiple definitions of influence radius in literature. In this work we define $r_{\mathrm{inf}}=G\,M_{\bullet}/\sigma^{2}_{\mathrm{inf}}$, where $\sigma_{\mathrm{inf}}$ is the local velocity dispersion. of $r_{\mathrm{inf}}\simeq 2.2\ \mathrm{pc}$ and the velocity dispersion at the influence radius corresponds to $\sigma_{\mathrm{inf}}=88\ \mathrm{km/s}$, computed according to the $M-\sigma$ by Gültekin et al. (2009). This system is used to compute the quantity $\left\rvert\Delta\mathcal{R}\right\rvert_{orb}$. The loss cone diffusivity $q=q(E)$ for the initially isotropic distribution and for the final profile (at $10$ Gyr) in this system is shown in Fig. 3. The net effect of mass segregation, which pushes stars to lower $E$ and stellar black holes (sBHs) towards higher $E$, is to increase $q$ for orbits with $a\simeq r_{c}(E)\lesssim 10^{-1}$ pc. In fact, in this region $m_{\star}\,\rho_{\star}\leq m_{\mathrm{sBH}}\,\rho_{\mathrm{sBH}}$ and therefore stellar mass black holes dominate two-body relaxation rates. In the rest of this work, we use the estimate of $q$ from the relaxed profile at $10$ Gyr.
@@
In Fig. 4 we show the predictions of our semi-analytical model for partial disruptions of a non-oscillating star. We consider both model A and model B stars; different values of the initial energy $E_{0}$ corresponding to $q=\{0.1,1.0,10.0\}$; different prescriptions for energy change in the core kick (Eqs. 10 and 11); and different models for tidal excitation (linear and non-linear).
@@
Irrespective of the specific models used for the computation of energy changes (i.e. tidal excitation and mass loss), there is a qualitative difference between the partial disruption of stars in the two stellar models driven by the different internal structure. The structure of model $B$ requires deeper penetration of the orbit (smaller $r_{\mathrm{p}}$) for the star to lose mass, so that the function $\delta m/m$ goes to zero at large pericentres faster than to stars described by model A. Moreover, model B corresponds to a larger energy stored into oscillations, with tidal oscillations being dominant at large $r_{\mathrm{p}}$. On the other hand, partial disruptions of model $A$ stars are dominated by the effects of mass loss at larger pericentres. In general, the linear expansion seems to perform well compared to non-linear estimates when mass loss is below approximately $5\%$. While in the case of model A Eq. 11 predicts a larger energy increase with respect to Eq. 10, the opposite is true in the case of model B given the larger pericentres involved in units of $r_{\mathrm{TDE}}$.
@@
To overview the full energy range, in Fig. 5 we show the qualitative effect of single passages pf a non-oscillating star at the pericentre as a function of the initial energy $E$ and angular momentum $J$ for model A and model B; we also report the radius $r_{c}$ of a circular orbit at energy $E$ for better interpretation. Here, we use the combination of energy change given by Eq. 11 and non-linear tidal excitations of Eq. 12. We plot the contours that correspond to $\Delta E/E=-100\%$, $|\Delta E/E|=10\%$ and $\delta m/m<1\%$, and define some (arbitrary) thresholds to classify the possible outcomes:
@@
  • Negligible change in the orbit, significant change in the mass. This happens when $|\Delta E|/E<10\%$ and $|\delta m|/m>1\%$ (aqua green).
  • Significant change in the orbit, significant change in the mass. This happens when $10\%<|\Delta E|/E<100\%$ and $|\delta m|/m>1\%$ (green).
@@
The level curves of model A are non-monotonic, and for $E\lesssim 3\,\sigma^{2}_{\mathrm{inf}}$ there are two disconnected curves corresponding to $\delta E=-10\%$. This is consistent with the trend of $E_{\mathrm{fin}}/E_{0}$ shown in Fig. 4: for model A, this ratio has a maximum (with $E_{\mathrm{fin}}/E_{0}>1$) and a minimum (with $E_{\mathrm{fin}}/E_{0}<1$ ). The structure of the levels $\delta E=\pm 10\%$, moving from small to large pericentres, is the following: (i) for any value of $E_{0}$ there is a small value of $r_{p}$ such that $E_{\mathrm{fin}}=0.9E_{0}$ and $\delta E=-10\%$; (ii) if the maximum of $E_{\mathrm{fin}}$ is larger than $1.1E_{0}$, there are two values of $r_{p}$ such that $E_{\mathrm{fin}}=1.1E_{0}$ and $\delta E=+10\%$; (iii) whenever the local minimum of $E_{\mathrm{fin}}$ is below $0.9\,E_{0}$, there are other two values of $r_{p}$ such that $E_{\mathrm{fin}}=0.9E_{0}$ and $\delta E=-10\%$. The level curves of model B have a simpler pattern, since $E_{\mathrm{fin}}/E_{0}$ only has a maximum with $E_{\mathrm{fin}}/E_{0}>1$ and is asymptotic from above to $1$ at large $r_{p}$. Therefore, the structure of the levels $\delta E=\pm 10\%$, moving from small to large pericentres, is the following: (i) for any value of $E_{0}$ there is a small value of $r_{p}$ such that $E_{\mathrm{fin}}=0.9E_{0}$ and $\delta E=-10\%$; (ii) if the maximum of $E_{\mathrm{fin}}$ is larger than $1.1E_{0}$, there are two values of $r_{p}$ such that $E_{\mathrm{fin}}=1.1E_{0}$ and $\delta E=+10\%$.
Overall, when $\delta m<1\%$ model A stars have smaller energy fluctuations, while the tidal excitation is strongly relevant in model B for semimajor axis (estimated through $r_{c}$) comparable to the influence radius. On the other hand, mass loss effects are dominant for model A and the total energy change is positive (i.e, the resulting orbit is more bound) only in an interval of values at small pericentres. Close to $r_{\mathrm{p}}/r_{\mathrm{TDE}}=\eta_{\mathrm{TTDE}}$, the energy fluctuation steeply decreases and becomes negative, so that the remnant is ejected in both models.
@@ 3.1 Full loss cone regime
When $q\gtrsim 1$, relaxation acts on timescales shorter than times between two partial disruption events. In this regime, single passages grazing or entering the loss cone radius will alter the orbital parameters once per period $P$, while the timescale of relaxation in angular momentum is $P/q$. This implies that only a small fraction of the stars instantaneously moving on an orbit with pericentre smaller than the tidal disruption radius will actually manage to reach such a pericentre, but they are likely to be scattered to a different (safe) orbit before reaching it. As a consequence, the probability of having a pTDE and staying on the same orbit to undergo a second disruption is small, and virtually negligible since $q$ grows quickly as the orbital energy $E$ decreases (see Fig. 3).
A fraction of partial disruptions of loosely bound stars may result in the ejection of the remnant at a pericentre larger than $r_{\mathrm{TDE}}$ (see Fig. 5), and this may be compatible with classical loss cone theory. In fact, since they both imply an instantaneous removal of the subject star from the distribution, ejections can be treated together with total disruptions in the standard approach to the loss cone, where particles are removed when they reach a pericentre below the critical value $r_{\mathrm{TDE}}$. Therefore, by identifying a suitable, larger threshold radius $r_{\mathrm{LC}}>r_{\mathrm{TDE}}$ that roughly distinguishes the region of ejections plus total disruptions, classical loss cone theory may be simply adapted to provide a good description of the system. Such a value of $r_{\mathrm{LC}}$ weakly depends on energy, but this does not affect the estimation of the corresponding Cohn-Kulsrud solution.
@@ 3.2 Empty loss cone regime
Figure 6: Effects of repeated partial disruptions with initial pericentre $r_{\mathrm{p}}$ (in units of $r_{\mathrm{TDE}}$) according to our simple step-by-step evolution for $q=0.1$ (upper figure), $q=0.01$ (middle figure) and $q=0.001$ (lower figure); the corresponding energy for our fiducial model is also reported in units of $\sigma^{2}_{\mathrm{inf}}$ and the orbital period $P$ in years. The first column refers to model A stars, the second column to model B stars. In the upper panel of each figure, we show the outcome of particles completing less than $N_{q}=1/q$ orbits, distinguishing between total disruptions ($r_{\mathrm{p}}<r_{\mathrm{TDE}}$) and ejections ($E_{\mathrm{fin}}<0$); we dub as survival the stars that complete $N_{q}$ orbits. In the middle panel, we report the ratio between the final and initial energy, showing the $90\%$ line (blue) and the $110\%$ line (orange) for reference. In the lower panel, we show the fraction of completed orbits $N_{\mathrm{orb}}/N_{q}$. We mark with a vertical dashed line the threshold pericentre, i.e. the smallest pericentre allowing the stars to complete $N_{q}$ orbits. Model A stars show a reduced orbital energy as they start closer to the threshold pericentre, meaning that they are less bound to the central MBH. On the other hand, model B stars show significant energy fluctuations slightly biased towards less bound orbits when computing the effect of asymmetric mass loss according to Gafton et al. (2015).
@@
For each pericentre passage, we will consider the mass lost according to Eq. (5) (model A) or Eq. (7) (model B) and the two prescriptions for the corresponding decrease in specific binding energy via core kicks, also including the exchange of energy between stellar oscillations and orbital motion according to Eq. (12). We remark that the perturbation due to tidal excitation is positive-definite (i.e. the orbit becomes more bound) only at the first pericentre passage. At subsequent passages, tidal deformation will inject or subtract energy to the stellar orbit depending on the phase of the ongoing oscillation (induced by the previous passage) when the star returns to pericentre. To include this effect, we use Eq. 12 to determine the amplitude of the energy exchanged but, from the second pericentre passage onwards, we give it a random sign at each passage.
@@
We treat $R_{\star}$ and $\eta_{\mathrm{TTDE}}$ as fixed parameters throughout the evolution, since modelling the readjustment of the star requires to directly solve the complex hydrodynamic evolution of the stellar structure [^6] Since we are considering orbital periods of at most $10^{4}$ yr, much less than the Kelvin-Helmholtz timescale ($\sim 1-100$ Myr; Kippenhahn & Weigert, 1990), dramatic mass loss and energy injection through the dissipation of nonlinear oscillation modes can in principle seriously affect the internal structure of the star.. In Appendix A we explore a non-fiducial scenario where the star can inflate due to nonlinear tides. In this case, part of the energy stored in oscillations is converted into internal energy through an enlargement of the radius. Finally, we remark that the procedure composed of steps $1-4$ implies that the specific angular momentum is constant throughout the repeated disruptions.
In Fig. 6 we show the results of our simulations by reporting the outcome (survival, total disruption or ejection), the final energy and the number of repeating partial disruption orbits for each simulation as a function of its initial pericentre $r_{\mathrm{p}}$. For both models and for any value of $q$, we see that repeating the partial disruption process can lead to total disruption and, in a minority of cases, to ejection at pericentres larger than $r_{\mathrm{TDE}}$ and $\eta_{\mathrm{TDE}}\,r_{\mathrm{TDE}}$. The threshold radius (reported in units of $r_{\mathrm{TDE}}$ in each sub-figure) increases with $N_{q}$ and therefore decreases with $q$. There are, however, some significant differences between the repeated partial disruption of a star described by model A and one described by model B.